OpenStudy (inowalst):

(x/100) + (y/66) = (3x+5y)/z. What is z?

OpenStudy (perl):

first i would multiply both sides by z*100*66, that will clear fractions

OpenStudy (inowalst):

So I would just be left with the variables?

OpenStudy (inowalst):

On the right side.

sammixboo (sammixboo):

I would say multiply all sides by z to get \(\dfrac{x}{100}+\dfrac{y}{66} = \dfrac{3x+5y}{z}\) \(\dfrac{x}{100}z+\dfrac{y}{66}z = \dfrac{3x+5y}{z}z\) \(\dfrac{1}{100}xz+\dfrac{1}{66}yz = 3x+5y\)

sammixboo (sammixboo):

Right?

OpenStudy (inowalst):

I'm a little confused..

OpenStudy (theraggedydoctor):

Hm.